40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 8779     Entry time: Fri Jun 28 02:12:21 2013
Author: manasa 
Type: Update 
Category: LSC 
Subject: PRMI + X arm ALS  

X arm stabilized using ALS while PRMI stayed locked

[Rana, Lisa, Jenne, Manasa]

Attachment 1

Time series : ALS enabled at t = 0 and disabled at t = 95s

PRMI_XALS_Jun28.png

What we did:
1. Jenne will elog about ASC (POP QPD) updates.
2. Found the beat note between Xarm green and PSL green.
3. Stabilized arm fluctuation by enabling ALS servo.
4. Scanned the arm for carrier resonance by ramping on the offset and set the offset such that we had IR resonating (TRX fluctuated between 0.1 and 0.8 counts).
5. Disabled the ALS servo and locked PRMI using AS55 for MICH and REFL33 for PRCL.
6. Enabled ALS.

Discussion:
Enabling ALS to detune the arm out of resonance kept PRMI locked  (currently for a span of few tens of seconds). However we could not see PRMI locked as stably compared to when the arms are misaligned. Everytime the offset was set IR to resonate, the PRMI was kicked out of lock.

Also there is some leakage at the arm transmission when PRMI was locked. The leakage was visible at ETMX transmission as flashes in different higher order modes indicating the not-so sufficient ALS stability. The leakage sets an offset at TRX measuring 0.01-0.05 counts.

To do list:
The ALS_OFFSETTER1 has to be calibrated in FSR. We were giving random offsets to do the offset scan.

Misc:
Installed a filter before ETMXT camera to remove the refl green. (Note to myself: The filter needs to go on a better mount/adapter).

ELOG V3.1.3-